# DevFD: Developmental Face Forgery Detection by Learning Shared and Orthogonal LoRA Subspaces

### v2

@@ Appendix B Details of Orthogonal Gradient
To estimate the gradient space of LoRA and orthogonalize it with the subspace of LoRA, we rely on a key proposition Liang and Li (2023) that the gradients of linear layers will lie within the input space. Here, we provide proof of this proposition. For a linear layer, let its weight matrix be denoted as $\mathbf{W}\in\mathbb{R}^{d_{I}\times d_{O}}$. Its forward process can be expressed as:
@@ C.1 Experimental Fairness
First, at the data level, we use the same amount of subtask data as the compared methods, i.e., 100 videos with 20 sampled frames per video, which is identical to methods such as DFIL Pan et al. (2023) and SUR-LID Cheng et al. (2024). Second, at the model level, we employ a backbone with a similar order of magnitude of parameters as the compared methods: ViT-B/16, thus avoiding the use of a larger model to ensure experimental fairness. It is worth noting that our trainable parameters consist only of the parameters of two LoRAs, which is approximately 1/10 of the trainable parameters of the compared methods. Third, in terms of training strategies, many methods such as DFIL Pan et al. (2023) and SUR-LID Cheng et al. (2024) utilize a Replay Set, which allows access to data from previous tasks—a practice may prohibited in many continual learning scenarios. The proposed DevFD method completely isolates any data between subtasks. This not only demonstrates that we achieve the best results under stricter constraints but also indicates the broader applicability of our method. Finally, the comparative metric data used are the best results extracted directly from the original papers. We achieve an undeniably leading position in absolute performance.
@@ C.3 More Details of Task10 Long Sequence Experiment
Existing continual learning methods for forgery detection typically employ task sequences with only four subtasks. We provide results for a continual learning task sequence with 10 subtasks based on DF40 Yan et al. (2024) and reproduce three high-performing continual learning methods for comparison: CoReD Kim et al. (2021), DFIL Pan et al. (2023), and HDP Sun et al. (2024). We start with the original task sequence [Hybrid, FR, FS, EFS], where Face-Swapping (FS): BlendFace, Face-Reenactment (FR): MCNet, and Entire Face Synthesis (EFS): StyleGAN3 are used as our initial learning sequence and expand it to 10 tasks. We carefully select the remaining six methods while maintaining diversity. Specifically, for Face-Swapping (FS), we choose SimSwap and InSwapper; for Face-Reenactment (FR), we select Wav2Lip and SadTalker; and for Entire Face Synthesis (EFS), we choose StyleGAN2 and StyleGAN-XL. We then randomly shuffle the last six methods to form the complete task sequence with 10 subtasks: [FFpp, MCNet, BlendFace, StyleGAN3, StyleGAN-XL, Wav2Lip, SadTalker, StyleGAN2, SimSwap, InSwapper]. Analysis reveals that EFS methods may have significant differences from FS and FR methods, leading to sharp increases in forgetting rates during transitions between learning EFS methods and FS & FR methods, such as 5->6 and 7->8.
C.4 Effect of orthogonal Subspace and Orthogonal Gradients
@@
C.5 Effect of ViT Backbone
We employ a Vision Transformer (ViT) pre-trained by CLIP Radford et al. (2021) as our initial backbone for training. To explore how different backbones affect model performance, we test various ViT models Ilharco et al. (2021) with varying sizes of parameters. The results of these experiments are detailed in Table 8.

### v3

@@ Appendix B Details of Orthogonal Gradient
To estimate the gradient space of LoRA and orthogonalize it with the subspace of LoRA, we rely on a key proposition api that the gradients of linear layers will lie within the input space. Here, we provide proof of this proposition. For a linear layer, let its weight matrix be denoted as $\mathbf{W}\in\mathbb{R}^{d_{I}\times d_{O}}$. Its forward process can be expressed as:
@@ C.1 Experimental Fairness
First, at the data level, we use the same amount of subtask data as the compared methods, i.e., 100 videos with 20 sampled frames per video, which is identical to methods such as DFIL dfil and SUR-LID surlid. Second, at the model level, we employ a backbone with a similar order of magnitude of parameters as the compared methods: ViT-B/16, thus avoiding the use of a larger model to ensure experimental fairness. It is worth noting that our trainable parameters consist only of the parameters of two LoRAs, which is approximately 1/10 of the trainable parameters of the compared methods. Third, in terms of training strategies, many methods such as DFIL dfil and SUR-LID surlid utilize a Replay Set, which allows access to data from previous tasks—a practice that is prohibited in many continual learning scenarios. The proposed DevFD method completely isolates any data between subtasks. This not only demonstrates that we achieve the best results under stricter constraints but also indicates the broader applicability of our method. Finally, the comparative metric data used are the best results extracted directly from the original papers. We achieve an undeniably leading position in absolute performance.
@@ C.3 More Details of Task10 Long Sequence Experiment
Existing continual learning methods for forgery detection typically employ task sequences with only four subtasks. We provide results for a continual learning task sequence with 10 subtasks based on DF40 df40 and reproduce three high-performing continual learning methods for comparison: CoReD cored, DFIL dfil, and HDP hdp. We start with the original task sequence [Hybrid, FR, FS, EFS], where Face-Swapping (FS): BlendFace, Face-Reenactment (FR): MCNet, and Entire Face Synthesis (EFS): StyleGAN3 are used as our initial learning sequence and expand it to 10 tasks. We carefully select the remaining six methods while maintaining diversity. Specifically, for Face-Swapping (FS), we choose SimSwap and InSwapper; for Face-Reenactment (FR), we select Wav2Lip and SadTalker; and for Entire Face Synthesis (EFS), we choose StyleGAN2 and StyleGAN-XL. We then randomly shuffle the last six methods to form the complete task sequence with 10 subtasks: [FFpp, MCNet, BlendFace, StyleGAN3, StyleGAN-XL, Wav2Lip, SadTalker, StyleGAN2, SimSwap, InSwapper]. Analysis reveals that EFS methods may have significant differences from FS and FR methods, leading to sharp increases in forgetting rates during transitions between learning EFS methods and FS & FR methods, such as 5->6 and 7->8.
Appendix D More Experimental Results
D.1 Parameter Analysis
D.2 Effect of orthogonal Subspace and Orthogonal Gradients
@@
D.3 Effect of ViT Backbone
We employ a Vision Transformer (ViT) pre-trained by CLIP clip as our initial backbone for training. To explore how different backbones affect model performance, we test various ViT models openclip with varying sizes of parameters. The results of these experiments are detailed in Table 8.